OpenStudy (anonymous):

Find the distance between the points (4,-3) and (-6,2) Find the distance between the points (4,-3) and (-6,2) @Mathematics

OpenStudy (anonymous):

use distance formula D=sqrt((x-xo)^2+(y-yo)^2)

OpenStudy (anonymous):

d^2=(x2-x1)^2 + (y2-y1)^2 d^2=(-6-4)^2 + (2--3)^2 d^2=100+25 d^2=125 d=sqrt125 d=11.18units

OpenStudy (anonymous):

sqrt 125